# Covariance scanning for adaptively optimal change point detection in high-dimensional linear models

### v1

@@ 5 Simulation study
We set $\bm{\Sigma}=\mathbf{I}$ with $p=900$, and $\bm{\beta}_{0}=-\bm{\beta}_{1}=\rho\bm{\delta}$ with $\rho\in\{1,2,4\}$, where $\bm{\delta}=(\delta_{1},\ldots,\delta_{p})^{\top}$ with $\delta_{i}=0$ for $i\notin\mathcal{S}\subseteq[p]$. The support $\mathcal{S}$ is sampled uniformly with cardinality $|\mathcal{S}|=\mathfrak{s}$, and $(\delta_{i},\,i\in\mathcal{S})$ is drawn uniformly from the unit sphere in $\mathbb{R}^{\mathfrak{s}}$. The sparsity level $\mathfrak{s}$ ranges over $\{1,2,3,4,7,11,18,30,49,79,129,210,341,554,900\}$, a subset of $[p]$ with values approximately evenly spaced on a log scale. We consider sample sizes $n\in\{300,600,900,1200\}$ with the change point located at $\theta=n/4$; see Figure 3 that expands on Figure 1. As noted in Section 2.1, we observe a clear phase transition around $\mathfrak{s}=\sqrt{p}$, which marks where the performance of the methods designed for handling the sparse regime (McScan, MOSEG and CHARCOAL), is overtaken by that of QcScan that performs uniformly across all $\mathfrak{s}$; OcScan and OcScan.R are able to enjoy the good performance of McScan and OcScan adaptively to the unknown sparsity, with the latter performing marginally better in the sparse regime where $\bm{\delta}$ itself is also sparse. Figure 4 additionally confirms that the good performance of the proposed covariance scanning methods is achieved with a fraction of time taken by MOSEG or CHARCOAL. OcScan.R involves a Lasso-type estimation of $\bm{\delta}$ and thus is relatively slower, but still is much faster than MOSEG that requires the computation of Lasso estimators of local regression parameters over time.
@@
The vector $\bm{\delta}_{0}$ itself has $\mathfrak{s}_{\bm{\delta}}$ non-zero elements taking values from $\{1,-1\}$ at random locations.
@@
The covariance weighted vector $\bm{\Sigma}^{1/2}\bm{\delta}_{0}$ has $\mathfrak{s}$ non-zero elements taking values from $\{1,-1\}$ at random locations.
The sparsity levels range over $\{1,2,4,7,14,27,53,103,200\}$, a subset of $[200]$ with values approximately evenly spaced on a log scale. In the special case of $\gamma=0$, the two types of sparsity coincide and, if further $\nu=0$, the scenario is similar to (M1) with the only difference in generating non-zero entries. We fix $n=300$ and $\theta=75$, and vary $p\in\{200,400\}$. See Figures 2, 5 and 6, and also those given in Appendix C.4. As discussed in Section 2.2, we observe that the inherent sparsity better captures the phase transition demonstrated by the performance of methods geared towards the sparse and dense regimes, respectively. Regardless of how the sparsity level is determined, OcScan is shown to adaptively enjoy the good performance of McScan or QcScan, depending on which method outperforms the other. The refinement step (Section 4.2) adopted by OcScan.R appears to be effective in the sparse regime (both in terms of $\mathfrak{s}_{\delta}$ and $\mathfrak{s}$) particularly as the size of change increases with $\rho$.
@@ B.1.1 Proof of Theorem 1
Let $\mathbf{1}^{\pm}_{S}$ denote a $p$-dimensional vector with its $i$-th entry being $1$ or $-1$ if $i\in S$ and being zero if $i\not\in S$, and introduce $\mathcal{U}=\bigl{\{}\bm{u}=({\kappa}/{\sqrt{\mathfrak{s}}})\mathbf{1}^{\pm}_{S}\in\mathbb{R}^{p}:S\subseteq[p],\;|S|=\mathfrak{s}\bigr{\}}$, whose cardinality is $2^{\mathfrak{s}}{p\choose\mathfrak{s}}$. Define
@@
By symmetry of Rademacher distribution, we can find an event $A$ such that $\mathsf{P}(A)=1/2$, $A\subseteq\{Z_{|S\cap[\mathfrak{s}]|}\geq 0\}$ and $A^{c}\subseteq\{Z_{|S\cap[\mathfrak{s}]|}\leq 0\}$. As $42\rho/(1-\rho^{2})\leq 8/15\leq 3/$ 4, we apply again Lemma B.1 and obtain
@@ B.2.3 Proof of Theorem 5
Consider some $m\in\mathcal{D}\cap(0,2^{-l^{*}-1}n)$, for which it holds that $42^{-l^{*}-1}n-m\geq 2^{-l^{*}-2}n\geq\theta/$ 4. Then, by Proposition 4 and Lemma B.12, we have for any $m\in\mathcal{D}\cap(0,2^{-l^{*}-1}n)$,
@@
Based on Step 1, provided that $\mathcal{A}_{1}$ holds, it is sufficient to consider the case where we have $(s,e)=(k^{*}/2,2k^{*})$ at the start of the OS algorithm. Besides, it is enough to consider the pairs of the probe points $(t,u)$ considered by the OS algorithm which lie on the same side of $\theta$, since otherwise no matter which side is dropped off, no mistake occurs. Thus we define
@@
such that t,u lie on the same side of θ with |$\quad\left.\text{such that $t,u$ lie on the same side of $\theta$with $|t-\theta|<|u-\theta|$},\right.$
@@ B.2.6 Supporting lemmas
Let $\mathbf{x}_{t}\sim_{\text{iid}}\mathcal{N}_{p}(\mathbf{0},\bm{\Sigma})$ for $t\in[n]$, $p\geq n$ and $\mathbf{X}=[\mathbf{x}_{1},\ldots,\mathbf{x}_{n}]^{\top}$. Consider some $0<k<n$. For all $z\geq 0$, any $\bm{\alpha},\bm{\beta}\in\mathbb{R}^{p}$ and some constant $C>0$, we have
@@
To proceed further, we need some notation. For positive integers $d$ and $q$, we denote a $d$-indexed matrix $\mathbf{C}=(c_{i_{1},\ldots,i_{d}})_{i_{1},\ldots,i_{d}=1}^{q}$; also we write $\mathbf{i}=(i_{1},\ldots,i_{d})\in[q]^{d}$ and $\mathbf{i}_{J}=(i_{l},\,l\in J)$ for $J\subseteq[d]$. Also, let $\mathcal{P}_{d}$ the set of partitions of $[d]$ into non-empty disjoint subsets. Given a partition $\mathcal{J}=\{J_{1},\ldots,J_{\ell}\}$, we define the tensor product norm:
@@
Denoting by $\mathbf{C}=[c_{(t_{1},i_{1}),(t_{2},i_{2}),(t_{3},i_{3}),(t_{4},i_{4})}]$ with $(t_{l},i_{l})\in[n]\times[p]$, we derive a bound on $\|\mathsf{E}(\partial^{4}f(\mathbf{Z}))\|_{\mathcal{J}}$ from that on $\|\mathbf{C}\|_{\mathcal{J}}$ by triangle inequality, as $[\mathsf{E}(\partial^{4}f(\mathbf{Z}))]_{\mathbf{i}_{1},\mathbf{i}_{2},\mathbf{i}_{3},\mathbf{i}_{4}}=\sum_{\pi}C_{\mathbf{i}_{\pi(1)},\mathbf{i}_{\pi(2)},\mathbf{i}_{\pi(3)},\mathbf{i}_{\pi(4)}}$, where the sum is taken over all permutations $\pi$ of $[4]$. When $\mathcal{J}=\{[4]\}$ (hence $|\mathcal{J}|=1$), it holds
@@ B.3.2 Proof of Proposition 9
$|\widehat{\theta}_{\text{\tiny R}}-\theta|$ $\lesssim|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{-2}\Psi^{2}\log(p\vee n)\max\left\{1,\,\frac{\|\bm{\Sigma}\|^{2}\Psi^{2}\mathfrak{s}_{\delta}^{2}\log(p\vee n)}{\underline{\sigma}^{2}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}\Delta}\right\}$

### v2

@@ 5 Simulation study
[…]
@@
We further consider the case where $\mathsf{Cov}(\mathbf{x}_{t})$ is rank deficient, which is accommodated by our theoretical framework but typically excluded under standard assumptions in the existing literature. For this, we retain the same data generation process as described above, with the modification pertaining to the construction of $\mathbf{x}_{t}$ only: We generate $\mathbf{x}_{t}=\mathbf{U}_{r}\widetilde{\mathbf{x}}_{t}$ for some $r\leq p$, where $\mathbf{U}_{r}\in\mathbb{R}^{p\times r}$ is a randomly generated matrix with orthonormal columns, and independently $\widetilde{\mathbf{x}}_{t}\sim_{\text{iid}}\mathcal{N}_{r}(\bm{0},\mathbf{I})$. It follows that $\mathsf{Cov}(\mathbf{x}_{t})=\mathbf{U}_{r}\mathbf{U}_{r}^{\top}$ which is of rank $r$. As illustrated in Figure 5, the proposed covariance scanning methods exhibit stable performance across varying $r\in\{p/4,p/2,p\}$ and their regime-specific behaviour is comparable to that observed when $\bm{\Sigma}=\mathbf{I}$. On the other hand, the performance of MOSEG and CHARCOAL grows worse as $r$ decreases, even when $\bm{\delta}$ is highly sparse and $n=1200>p=900$. Additional results are presented in Section C.4.
@@
Figure 5: Estimation performance of MOSEG, CHARCOAL, McScan, QcScan, OcScan and OcScan.R in modified (M1) where $n=1200$, $p=900$ and $\bm{\Sigma}=\mathbf{U}_{r}\mathbf{U}_{r}^{\top}$ with varying $r\in\{p/4,p/2,p\}$, where $\mathbf{U}_{r}\in\mathbb{R}^{p\times r}$ is randomly generated with orthonormal columns. The $x$-axis denotes $\mathfrak{s}_{\delta}=|\bm{\delta}|_{0}$. In each scenario, the results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The vertical dashed lines mark where $\mathfrak{s}_{\delta}=\sqrt{p}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
@@
The vector $\bm{\delta}_{0}$ itself has $\mathfrak{s}_{\bm{\delta}}$ non-zero elements taking values from {$,-1\}$ 1 } at random locations.
@@
The covariance weighted vector $\bm{\Sigma}^{1/2}\bm{\delta}_{0}$ has $\mathfrak{s}$ non-zero elements taking values from {$,-1\}$ 1 } at random locations.
[…]
@@ B.1.1 Proof of Theorem 1
Let $\mathbf{1}^{\pm}_{S}$ denote a $p$-dimensional vector with its $i$-th entry being 111 or 1-1- 1 if $i\in S$ and being zero if $i\not\in S$, and introduce $\mathcal{U}=\bigl{\{}\bm{u}=({\kappa}/{\sqrt{\mathfrak{s}}})\mathbf{1}^{\pm}_{S}\in\mathbb{R}^{p}:S\subseteq[p],\;|S|=\mathfrak{s}\bigr{\}}$, whose cardinality is $2^{\mathfrak{s}}{p\choose\mathfrak{s}}$. Define
@@
By symmetry of Rademacher distribution, we can find an event $A$ such that $\mathsf{P}(A)=1/2$, $A\subseteq\{Z_{|S\cap[\mathfrak{s}]|}\geq 0\}$ and $A^{c}\subseteq\{Z_{|S\cap[\mathfrak{s}]|}\leq 0\}$. As $2\rho/(1-\rho^{2})\leq 8/15\leq 3/4$, we apply again Lemma B.1 and obtain
@@ B.2.3 Proof of Theorem 5
Consider some $m\in\mathcal{D}\cap(0,2^{-l^{*}-1}n)$, for which it holds that $2^{-l^{*}-1}n-m\geq 2^{-l^{*}-2}n\geq\theta/4$. Then, by Proposition 4 and Lemma B.12, we have for any $m\in\mathcal{D}\cap(0,2^{-l^{*}-1}n)$,
@@
[…]
$\quad\left.\text{such that $t,u$ lie on the same side of $\theta$ with $|t-\theta|<|u-\theta|$},\right.$
@@ B.2.6 Supporting lemmas
[…]
@@
[…]
@@
Denoting by $\mathbf{C}=[c_{(t_{1},i_{1}),(t_{2},i_{2}),(t_{3},i_{3}),(t_{4},i_{4})}]$ with $(t_{l},i_{l})\in[n]\times[p]$, we derive a bound on $\|\mathsf{E}(\partial^{4}f(\mathbf{Z}))\|_{\mathcal{J}}$ from that on $\|\mathbf{C}\|_{\mathcal{J}}$ by triangle inequality, as $[\mathsf{E}(\partial^{4}f(\mathbf{Z}))]_{\mathbf{i}_{1},\mathbf{i}_{2},\mathbf{i}_{3},\mathbf{i}_{4}}=\sum_{\pi}C_{\mathbf{i}_{\pi(1)},\mathbf{i}_{\pi(2)},\mathbf{i}_{\pi(3)},\mathbf{i}_{\pi(4)}}$, where the sum is taken over all permutations $\pi$ of [4][4][ 4 ]. When $\mathcal{J}=\{[4]\}$ (hence $|\mathcal{J}|=1$), it holds
@@ B.3.2 Proof of Proposition 9
$|\displaystyle|\widehat{\theta}_{\text{\tiny R}}-\theta|$$\lesssim|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{-2}\Psi^{2}\log(p\vee n)\max\left\{1,\,\frac{\|\bm{\Sigma}\|^{2}\Psi^{2}\mathfrak{s}_{\delta}^{2}\log(p\vee n)}{\underline{\sigma}^{2}|\bm{\Sigma}^{1/2}\bm{\delta}|_{2}^{2}\Delta}\right\}$
@@ C.4 Additional simulations results under (M1)
Additional simulation results under the rank deficient variant of scenario (M1) in Section 5 are given in Figures C.10 and C.11. Together with Figure 3, these figures demonstrate that the performance advantage of the proposed covariance scanning methods over the competitors (MOSEG and CHARCOAL) becomes more pronounced as the rank $r$ decreases.
Figure C.10: Estimation performance of MOSEG, CHARCOAL, McScan, QcScan, OcScan and OcScan.R in modified (M1) with rank deficiency $r=p/2$; CHARCOAL is only applicable in the last column where $n=1200>p=900$. The $x$-axis denotes $\mathfrak{s}_{\delta}=|\bm{\delta}|_{0}$. In each scenario, the results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The vertical dashed lines mark where $\mathfrak{s}_{\delta}=\sqrt{p}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
Figure C.11: Estimation performance of MOSEG, CHARCOAL, McScan, QcScan, OcScan and OcScan.R in modified (M1) with rank deficiency $r=p/4$; CHARCOAL is only applicable in the last column where $n=1200>p=900$. The $x$-axis denotes $\mathfrak{s}_{\delta}=|\bm{\delta}|_{0}$. In each scenario, the results are based on 1000 repetitions, with median error curves shown alongside shaded regions representing the interquartile range. The vertical dashed lines mark where $\mathfrak{s}_{\delta}=\sqrt{p}$. The $x$-axis is shown on a log scale, and the $y$-axis on a squared root scale.
C.5 Additional simulations results under (M2)